# ELA-ZSON: Efficient Layout-Aware Zero-Shot Object Navigation Agent with Hierarchical Planning

### v1

@@ Abstract.
We introduce ELA-ZSON, an efficient layout-aware zero-shot object navigation (ZSON) approach designed for complex multi-room indoor environments. By planning hierarchically leveraging a global topologigal map with layout information and local imperative approach with detailed scene representation memory, ELA-ZSON achieves both efficient and effective navigation. The process is managed by an LLM-powered agent, ensuring seamless effective planning and navigation, without the need for human interaction, complex rewards, or costly training. Our experimental results on the MP3D benchmark achieves 85% object navigation success rate (SR) and 79% success rate weighted by path length (SPL) (over 40% point improvement in SR and 60% improvement in SPL compared to exsisting methods). Furthermore, we validate the robustness of our approach through virtual agent and real-world robotic deployment, showcasing its capability in practical scenarios. See https://anonymous.4open.science/r/ELA-ZSON-C67E/ for details.
Hierarchical Planning, Object Navigation, LLM Agent, Robotic Planning, Scene Understanding.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Robotic planningccs: Computing methodologies Mobile agentsccs: Computing methodologies Intelligent agentsccs: Computing methodologies Scene understandingccs: Computing methodologies Vision for roboticsccs: Computing methodologies Hierarchical representations
1. Introduction
@@
Advancements in Large Foundation Models (LFMs) and robotics research have brought household assistant robots closer to real-world deployment. A key capability for such robots is to locate a target and navigate to it based on user input (Zhu et al., 2017). Researchers have made significant progress in integrating Vision-Language Models (VLMs) to align user instructions with robot observations for prompted navigation and target localization using scene memory collected during exploration (Shah et al., 2023; Zhou et al., 2023; Dorbala et al., 2022; Chen et al., 2019; Vasudevan et al., 2021; Krantz et al., 2023). While complementary scene understanding and accurate instruction interpretation are essential for robotic navigation, two mainstream approaches have emerged: process-prompted and goal-oriented navigation.
@@
Recent advancements in Vision-Language Navigation (VLN) have focused on understanding and executing user instructions by leveraging Vision-Language Models (VLMs). These models enable robots to follow process-prompted instructions, allowing them to explore and navigate in unseen environments (Anderson et al., 2018b; Chen et al., 2019). However, these benchmarks are based on sparse panoramic observations, assuming known topologies, oracle navigation, and perfect localization, which do not reflect real-world deployment challenges (Krantz et al., 2020). VLN-CE has attempted to bridge this gap by removing unrealistic assumptions, translating observations into low-level controls in continuous environments (Krantz et al., 2020). The grounding of process-prompted instructions is in a step-by-step form, such as “Walk out of the bedroom. Turn right and walk down the hallway. At the end of the hallway turn left.”.
@@
Figure 1. Our proposed ELA-ZSON. An efficient ZSON approach with LLM-powered agent that realizes hierarchically planning based on topology and dense scene memory. The entire process is conducted automatically without costly training or complex rewards.
[…]
An efficient ZSON approach with LLM-powered agent that realizes hierarchically planning based on topology and dense scene memory. The entire process is conducted automatically without costly training or complex rewards.
@@
Figure 2. Overview of our proposed method. The LLM agent takes user instructions as input and manages the optional action choices according to the prompts and data flow. Optional actions include exploring and recording the scene, constructing scene memory representation, planning, and executing navigation. Obstacles and error recognition together with iterative attempts are available.
[…]
The LLM agent takes user instructions as input and manages the optional action choices according to the prompts and data flow. Optional actions include exploring and recording the scene, constructing scene memory representation, planning, and executing navigation. Obstacles and error recognition together with iterative attempts are available.
@@
While following step-by-step instructions is a necessary skill for household assistants, the goal-oriented task provides a simple optional interaction pattern for users. For instance, robotics should interpret simple instructions like “Go to the couch in the living room” or a picture of the desired couch. Since ZSON (Majumdar et al., 2022) proposed the ability to locate and navigate to the object of interest in a scene based on a user inquiry, without training on this specific setting, researchers have made significant efforts to improve retrieval processes, instance recognition abilities, and scene understanding (Guan et al., 2024; Sun et al., 2024; Zhou et al., 2023; Ramakrishnan et al., 2022; Yu et al., 2023). Despite these advancements, navigation in open scenes remains a challenging problem due to practical constraints that current algorithms have not fully addressed. Specifically, the following challenges persist: 1) While targets may not be directly observed, planning globally efficient paths based on scene memory without unnecessary detours is important. 2) Adapting to local scene changes that may occur compared to the initial scene memory. 3) Realizing effective planning without human aid, complex rewards, or costly training.
@@
Recent works in robotic planning have demonstrated diverse approaches. Works like L3MVN (Yu et al., 2023), ESC (Zhou et al., 2023), and VLFM (Yokoyama et al., 2024) employ LFMs to respectively map the scene with language-guided frontier exploration, probabilistic commonsense constraints, and visual-language similarity scoring, bridging semantic understanding with geometric planning. Taking global layouts into consideration, HOVSG (Werby et al., 2024) constructs open-vocabulary 3D scene topological graphs through feature clustering to enable navigation. However, the static topological way-points extracted from occupancy cannot ensure efficient path planning or manage unexpected local scene changes. To solve this dual-challenge, we propose leveraging a hierarchical framework uniquely integrates lightweight global topological map for coarse route planning with a local imperative approach that dynamically adjusts trajectories.
@@
In this work, we introduce ELA-ZSON, an efficient layout-aware zero-shot object navigation framework designed for complex multi-room indoor environments. ELA-ZSON leverages both topological layout relationships and detailed scene representations to achieve efficient navigation in a hierarchical framework. By employing the dual-level information, the planning process first queries the open-world semantics based on user instructions to generate a global navigable topometric path at the layout level. The global waypoints are then sequentially projected into the robot’s ego-centric observation space, enabling the robot to plan a dense local waypoint set that accounts for scenario changes. An LLM-powered agent is employed that operates without human interaction, complex reward design, or costly training on specific settings. The method requires only RGB-D observations as input and outputs navigation results as 3D waypoint sets, which are compatible with general platforms. The primary functions are shown in Fig.1. Evaluated on the widely-used MP3D dataset (Ramakrishnan et al., 2021), ELA-ZSON demonstrates superior navigation success rates (SR) and efficiency, as measured by success weighted by path length (SPL), compared to state-of-the-art (SOTA) methods. Additionally, we validate our framework through both virtual agent and real-world robotics, demonstrating its applicability, efficiency and robustness against unexpected scene changes.
@@
  • We present ELA-ZSON, an efficient layout-aware zero-shot object navigation approach that leverages hierarchical planning, managed by an LLM-powered agent without further rewards or training.
@@
  • Complex experiments are conducted on the MP3D dataset demonstrating a 85% SR and 79% SPL in object navigation. Deployments on both virtual agent and real robotics are provided to verify the applicability and robustness.
2. Related Works
@@
Figure 3. Hierarchical planning example. Global planning is conducted on the topological graph and generates $V=\{v_{1},v_{2},\dots,v_{n}\}$. Local planning, realized in IL approach, generates $W_{i,i+1}=\{w_{i}^{1},w_{i}^{2},\dots,w_{i}^{m}\}$ and detailed executable points.
[…]
Global planning is conducted on the topological graph and generates $V=\{v_{1},v_{2},\dots,v_{n}\}$. Local planning, realized in IL approach, generates $W_{i,i+1}=\{w_{i}^{1},w_{i}^{2},\dots,w_{i}^{m}\}$ and detailed executable points.
2.1. Process-prompted Vision and Language Navigation
@@
Vision-Language navigation requires agents to follow free-form linguistic instructions to navigate in unseen environments (Anderson et al., 2018b), necessitating an understanding of complex observations and the ability to interpret instructions for effective navigation. Pioneering works were conducted under the simplified assumption that environments could be abstracted as discrete connectivity graphs (Anderson et al., 2018b; Krantz et al., 2020). To bridge the gap for more realistic modeling, researchers have dedicated efforts to address the limitations imposed by task settings (Krantz et al., 2020; Zhang et al., 2024; Krantz et al., 2021; Xu et al., 2023; Anderson et al., 2021). However, the assumption of closed-set scene priors limits the agent to pre-mapped environments. Additionally, VLN evaluation metrics often emphasize strict path fidelity over goal discovery efficiency. On the contrary, we follow a goal-oriented paradigm and propose a solution involving open-set scene-specific representations built during exploration and zero-shot object navigation, executed in a human-interaction-free manner using an LLM-powered agent.
2.2. Goal-oriented Navigation
@@
Goal-oriented navigation involves interpreting multi-modal target descriptions, localizing, and navigating to the specified targets. Notable works (Yang et al., 2018; Du et al., 2020; Campari et al., 2020; Liang et al., 2021; Chaplot et al., 2020) have demonstrated impressive capabilities using large-scale reinforcement learning. However, these methods typically require substantial training data and computational resources and may suffer when generalizing across diverse environments and embodiments (Ramakrishnan et al., 2022). ZSON (Majumdar et al., 2022) trains navigation agents in target environments by mapping object goals to image-goal embeddings, emphasizing the ability to locate and navigate to objects based on inquiries without specific training in that setting. Further advancements have been made (Guan et al., 2024; Sun et al., 2024) to enhance image retrieval and semantic understanding, achieving higher success rates. Additionally, L3MVN (Yu et al., 2023) has improved exploration efficiency by constructing environment maps using frontier-based methods and leveraging LLM inference. ESC (Zhou et al., 2023) introduced soft commonsense constraints through probabilistic logic, enabling training-free zero-shot navigation by embedding object-room relationships, while VLFM (Yokoyama et al., 2024) eliminated text dependency by directly grounding visual features with language models through RGB-based similarity scoring, significantly accelerating semantic processing. Nonetheless, the absence of explicit global layout-level guidance limits their ability to ensure efficient path planning. Works like HOVSG (Werby et al., 2024) construct open-vocabulary 3D scene topological graphs through feature clustering to enable navigation. However, the static topological way-points extracted from occupancy cannot ensure efficient path planning or manage unexpected local scene changes.
3. Overview
@@
We propose a hierarchical robotic navigation framework for operation in unexplored indoor environments, guided by natural language instructions $T$, images $I$, or 3D locations $P$. The framework overview is shown in Fig.2. The planning process is conducted on a dual-level map with global topology and local dense representation built with collected RGB-D sequences ${i,d}$, where $i$ and $d$ denote RGB and depth observations, respectively. The scene representation consists of (1) an implicit neural function $F:\mathbb{R}^{3}\to\mathbb{R}^{n}$ maps 3D positions to vision-language embeddings; and (2) a topometric map $G=(V,E)$, where $V$ comprises region vertices $\mathbf{v}_{r}$ and entrance vertices connecting regions $\mathbf{v}_{e}$ and edges $E$ connecting these vertices.
@@
During planning, inputs $T/I$ are encoded into embeddings $\mathbf{E}$ using a pre-trained vision-language model. These embeddings are then queried in $F$ to identify the target position $p_{\text{end}}$ with the highest matching similarity. A global navigable path is computed between the robot’s current pose $p_{\text{start}}$ and $p_{\text{end}}$ using the topometric map $G$, generating a sparse sequence of waypoint vertices $V=\{v_{1},v_{2},\dots,v_{n}\}$. These waypoints are progressively refined during navigation in an imperative approach, with $W_{i,i+1}=\{w_{i}^{1},w_{i}^{2},\dots,w_{i}^{m}\}$ denoting the densified local waypoints between $v_{i}$ and $v_{i+1}$. Navigation proceeds incrementally: as the robot approaches $v_{i+1}$, the next segment $W_{i+1,i+2}$ is planned. The framework incorporates an LLM-powered agent designed to integrate essential data streams for effective task execution.
4. Method
4.1. Hierarchical Scene Representation
@@
Recent works such as HOV-SG (Werby et al., 2024), CLIO (Maggio et al., 2024), and Topo-Field (Hou et al., 2024) have demonstrated progress in representing scenes as topo-maps while retaining detailed content. Although topometric maps offer computational efficiency for downstream planning and navigation tasks, high-fidelity scene content from neural implicit representations enables semantic cue queries (Hou et al., 2024). Our approach represents 3D scenes combining a high-level topometric map and a detailed neural implicit representation.
@@
Given posed RGB-D sequences $\{i,d\}$, we train a neural implicit function $F$ (Hou et al., 2024) to match 3D points $p$ derived from depth and camera parameters with the encoded vision-language embedding $\mathcal{E}_{p}$ (Radford et al., 2021; Reimers and Gurevych, 2019). A topometric map $G=(V,E)$ is then extracted with vertices $V$ including region vertices $\mathbf{v}_{r}$ and entrance vertices $\mathbf{v}_{e}$ and edges $E$ connecting them. The subsequent scene representation can be denoted as:
(1) $F$ $:\mathbb{R}^{3}\to\mathbb{R}^{n},$
@@
Details of neural field and topology construction pipeline can be referred to in (Hou et al., 2024).
4.2. Hierarchical Planning
@@
4.2.1. Global Topology Plan
To ensure globally efficient path planning for inter-room navigation, our approach utilizes a topological graph to generate a sparse waypoint set that the robotic must traverse to reach the destination, typically located at room entrances or connecting doors. The process begins by unifying multi-modal instructions (images, text, or 3D positions) into a destination 3D position. The text/image instruction $T/I$ is first encoded using CLIP (Radford et al., 2021) and Sentence-BERT (Reimers and Gurevych, 2019) to generate the target embedding $\{C,S\}_{T/I}$. A point set $P^{*}$ is then sampled uniformly from the environment, and $F(P^{*})=\{(f_{v},f_{s})\}_{P^{*}}$, representing the vision and semantic embeddings of the corresponding 3D positions, is calculated. The cosine similarity between $\{C,S\}_{T/I}$ and $F(P^{*})$ determines the position $p_{\text{end}}\in P^{*}$ with the highest similarity to the instruction query (Hou et al., 2024). The topological graph $G$ is subsequently queried using the robot’s current position $p_{\text{start}}$ and $p_{\text{end}}$ to identify the shortest path connecting them. This results in a sequence of vertices $V=\{v_{1},v_{2},\dots,v_{n}\}$, where each $v_{i}$ represents a structural element, such as doors or corridors, that the robot must navigate through to reach the destination.
4.2.2. Local Ego-centric Plan
In the initial global planning phase, a sparse set of waypoints denoted as $V=\{v_{1},v_{2},\dots,v_{n}\}$, is established. These waypoints guide the robot’s inter-room navigation, ensuring efficient travel between distinct regions. To achieve intra-room movement and local navigation precision, dense waypoints, $W_{i,i+1}=\{w_{i}^{1},w_{i}^{2},\dots,w_{i}^{m}\}$, are computed between consecutive global waypoints $v_{i}$ and $v_{i+1}$. We realize local planning in an Imperative Learning (IL) approach with the target way-point $v_{i+1}$ in the global coordinates and current depth observation $d^{\prime}$ with pose $p^{\prime}$. First, $v_{i+1}$ is projected to the depth frame with the pose $p^{\prime}$. The depth input, encoded by a convolutional network (LeCun et al., 1995) into a depth embedding $\mathcal{O_{d^{\prime}}}$, together with the way-point feature $\mathcal{P^{\prime}}$ of $p^{\prime}$ is input into a planning network (Yang, 2023) that generates a set of local key points $W_{i,i+1}$ from
(2) $\tau=\mathcal{F}(\mathcal{O}_{d^{\prime}},\mathcal{P^{\prime}}),$
@@ 4.2.2. Local Ego-centric Plan
(3) $p_{\text{start}}\to p_{\text{end}}$ $:V=\{v_{1},v_{2},\dots,v_{n}\},$
@@
4.3. LLM-powered Navigation Agent
Figure 4. Data flow example of LLM-powered Agent. Instructed by the user input, the agent conducts prompts, applies optional actions, and gets feedback from the Robo-env interaction.
[…]
Instructed by the user input, the agent conducts prompts, applies optional actions, and gets feedback from the Robo-env interaction.
@@ 4.3. LLM-powered Navigation Agent
CoW (2022) (Gadre et al., 2022) 0.061 0.039 0.018 0.011 - -
ZSON (2022) (Majumdar et al., 2022) 0.255 0.126 0.106 0.049 0.146 0.073
L3MVN (2023) (Yu et al., 2023) 0.352 0.165 - - - -
ESC (2023) (Zhou et al., 2023) 0.392 0.223 0.065 0.037 - -
VLFM (2024) (Yokoyama et al., 2024) 0.364 0.175 - - - -
PixNav (2024) (Cai et al., 2024) 0.379 0.205 - - - -
HOV-SG (2024) (Werby et al., 2024) 0.404 0.236 - - - -
PSL (2024) (Sun et al., 2024) 0.424 0.192 0.165 0.075 0.230 0.114
ELA-ZSON(Ours) 0.856 0.797 0.786 0.701 0.673 0.582
Table 1. Quantitative comparison of object navigation and instance navigation tasks on MP3D dataset. For object navigation, text instruction is used as input. For instance navigation, text-prompted and image-prompted instructions are separately validated. The SR and SPL are employed as metrics.
@@
Table 2. Quantitative object navigation evaluation in real-world scene. Each navigation task is evaluated 10 times from a similar starting position to the same target. Obstacles means the number of obstacles we set after robotic exploration where robotics have to bypass to approach targets. Local/Global-Plan Retries means the number of replanning called by the agent during the 10 times of navigation. Path Length is the average path length robotics travel in 10 times navigation. Time cost is the average cost from starting points to destinations, where failure cases are not counted. SR is the navigation success rate.
@@
(4) $\{\mathbf{B},\mathbf{I},\mathbf{M},\mathbf{P},\mathbf{T},\mathbf{S},\mathbf{O},\mathbf{A}\}$
where $\mathbf{B}$ represents the task background information, structured as “You are an embodied robotic …”. $\mathbf{I}$ denotes the user’s instruction, which can be text, an image, or a 3D position. $\mathbf{M}$ is the hierarchical map of the environment, comprising a learned implicit function $F$ and a topological graph $G$ in JSON format. $\mathbf{P}$ includes the robot’s historical planning data. $\mathbf{T}$ refers to the robotic travel trajectory. $\mathbf{S}$ indicates the robot’s current status, such as its location and whether the planned path remains viable. $\mathbf{O}$ represents any optional skills the robot may use. $\mathbf{A}$ is the action decision made by the agent.
@@
  • Planning. Based on the scene representation $(F,G)$ and user instructions, the agent conducts planning described in Section 4.2 to fulfill the task.

### v2

# LOG-Nav: Efficient Layout-Aware Object-Goal Navigation with Hierarchical Planning

@@ Abstract
We introduce LOG-Nav, an efficient layout-aware object-goal navigation approach designed for complex multi-room indoor environments. By planning hierarchically leveraging a global topologigal map with layout information and local imperative approach with detailed scene representation memory, LOG-Nav achieves both efficient and effective navigation. The process is managed by an LLM-powered agent, ensuring seamless effective planning and navigation, without the need for human interaction, complex rewards, or costly training. Our experimental results on the MP3D benchmark achieves 85% object navigation success rate (SR) and 79% success rate weighted by path length (SPL) (over 40% point improvement in SR and 60% improvement in SPL compared to exsisting methods). Furthermore, we validate the robustness of our approach through virtual agent and real-world robotic deployment, showcasing its capability in practical scenarios.
Introduction
@@
Advancements in Large Foundation Models (LFMs) and robotics research have brought household assistant robots closer to real-world deployment. A key capability for such robots is to locate a target and navigate to it based on user input (zhu2017target). Researchers have made significant progress in integrating Vision-Language Models (VLMs) to align user instructions with robot observations for prompted navigation and target localization using scene memory collected during exploration (shah2023lm; zhou2023esc; dorbala2022clip; chen2019touchdown; vasudevan2021talk2nav; krantz2023iterative). While complementary scene understanding and accurate instruction interpretation are essential for robotic navigation, two mainstream approaches have emerged: process-prompted and goal-oriented navigation.
@@
Recent advancements in robitc navigation have focused on understanding and executing user instructions by leveraging Vision-Language Models (VLMs). These models enable robots to follow process-prompted instructions, allowing them to explore and navigate in unseen environments (anderson2018vln; chen2019touchdown). However, these benchmarks are based on sparse panoramic observations, assuming known topologies, oracle navigation, and perfect localization, which do not reflect real-world deployment challenges (vlnce). VLN-CE has attempted to bridge this gap by removing unrealistic assumptions, translating observations into low-level controls in continuous environments (vlnce). The grounding of process-prompted instructions is in a step-by-step form, such as “Walk out of the bedroom. Turn right and walk down the hallway. At the end of the hallway turn left.”.
@@
Figure 1: Our proposed LOG-Nav. An efficient object-goal navigation approach with LLM-powered agent that realizes hierarchically planning based on topology and dense scene memory. The entire process is conducted automatically without costly training or complex rewards.
@@
Figure 2: Overview of our proposed method. The LLM agent takes user instructions as input and manages the optional action choices according to the prompts and data flow. Optional actions include exploring and recording the scene, constructing scene memory representation, planning, and executing navigation. Obstacles, error recognition, and iterative attempts are available.
@@
While following step-by-step instructions is a necessary skill for household assistants, the goal-oriented task provides a simple optional interaction pattern for users. For instance, robotics should interpret simple instructions like “Go to the couch in the living room” or a picture of the desired couch. Since ZSON (majumdar2022zson) proposed the ability to locate and navigate to the object of interest in a scene based on a user inquiry, without training on this specific setting, researchers have made significant efforts to improve retrieval processes, instance recognition abilities, and scene understanding (guan2024loczson; psl; zhou2023esc; ramakrishnan2022poni; yu2023l3mvn). Despite these advancements, navigation in open scenes remains a challenging problem due to practical constraints. Specifically, the following challenges persist: 1) While targets may not be directly observed, planning globally efficient paths based on scene memory without unnecessary detours is important. 2) Adapting to local scene changes that may occur compared to the initial scene memory. 3) Realizing effective planning without human aid, complex rewards, or costly training.
@@
Recent works in robotic planning have demonstrated diverse approaches. Works like L3MVN (yu2023l3mvn), ESC (zhou2023esc), and VLFM (yokoyama2024vlfm) employ LFMs to respectively map the scene with language-guided frontier exploration, probabilistic commonsense constraints, and visual-language similarity scoring, bridging semantic understanding with geometric planning. Taking global layouts into consideration, HOVSG (werby23hovsg) constructs open-vocabulary 3D scene topological graphs through feature clustering to enable navigation. However, the static topological way-points extracted from occupancy cannot ensure efficient path planning or manage unexpected local scene changes. To solve this dual-challenge, we propose leveraging a hierarchical framework uniquely integrates lightweight global topological map for coarse route planning with a local imperative approach that dynamically adjusts trajectories.
@@
In this work, we introduce LOG-Nav, an efficient layout-aware object-goal navigation framework designed for complex multi-room indoor environments. LOG-Nav leverages both topological layout relationships and detailed scene representations to achieve efficient navigation in a hierarchical framework. By employing the dual-level information, the planning process first queries the open-world semantics based on user instructions to generate a global navigable topometric path at the layout level. The global waypoints are then sequentially projected into the robot’s ego-centric observation space, enabling the robot to plan a dense local waypoint set that accounts for scenario changes. An LLM-powered agent is employed that operates without human interaction, complex reward design, or costly training on specific settings. The method requires only RGB-D observations as input and outputs navigation results as 3D waypoint sets, which are compatible with general platforms. The primary functions are shown in Fig.1. Evaluated on the widely-used MP3D dataset (ramakrishnan2021hm3d), LOG-Nav demonstrates superior navigation success rates (SR) and efficiency, as measured by success weighted by path length (SPL), compared to exisiting methods. Additionally, we validate our framework through both virtual agent and real-world robotics, demonstrating its applicability, efficiency and robustness against unexpected scene changes.
@@
  • We present LOG-Nav, an efficient layout-aware object-goal navigation approach that leverages hierarchical planning, managed by an LLM-powered agent without further rewards or training.
@@
  • Complex experiments are conducted on the MP3D dataset demonstrating a 85% SR and 79% SPL in object navigation. Deployments on both virtual agent and real robotics are provided to verify the applicability.
Related Works
@@
Figure 3: Hierarchical planning example. Global planning is conducted on the topological graph and generates $V=\{v_{1},v_{2},\dots,v_{n}\}$. Local planning, realized in IL approach, generates $W_{i,i+1}=\{w_{i}^{1},w_{i}^{2},\dots,w_{i}^{m}\}$.
Process-prompted Navigation
@@
Vision-Language navigation requires agents to follow free-form linguistic instructions to navigate in unseen environments (anderson2018vln), necessitating an understanding of complex observations and the ability to interpret instructions for effective navigation. Pioneering works were conducted under the simplified assumption that environments could be abstracted as discrete connectivity graphs (anderson2018vln; vlnce). To bridge the gap for more realistic modeling, researchers have dedicated efforts to address the limitations imposed by task settings (vlnce; zhang2024navid; krantz2021waypoint; xu2023vision; anderson2021sim). However, the assumption of closed-set scene priors limits the agent to pre-mapped environments. Additionally, VLN evaluation metrics often emphasize strict path fidelity over goal discovery efficiency. On the contrary, we follow a goal-oriented paradigm and propose open-set scene-specific representations built during exploration and object navigation, executed in a human-interaction-free manner with LLM agent.
Goal-oriented Navigation
@@
Goal-oriented navigation involves interpreting multi-modal target descriptions, localizing, and navigating to the specified targets. Notable works (yang2018visual; du2020learning; campari2020exploiting; liang2021sscnav; chaplot2020object) have demonstrated impressive capabilities using large-scale reinforcement learning. However, these methods typically require substantial training data and computational resources and may suffer when generalizing across diverse environments and embodiments (ramakrishnan2022poni). ZSON (majumdar2022zson) trains navigation agents in target environments by mapping object goals to image-goal embeddings, emphasizing the ability to locate and navigate to objects based on inquiries without specific training in that setting. Further advancements have been made (guan2024loczson; psl) to enhance image retrieval and semantic understanding, achieving higher success rates. Additionally, L3MVN (yu2023l3mvn) has improved exploration efficiency by constructing environment maps using frontier-based methods and leveraging LLM inference. ESC (zhou2023esc) introduced soft commonsense constraints through probabilistic logic, enabling training-free zero-shot navigation by embedding object-room relationships, while VLFM (yokoyama2024vlfm) eliminated text dependency by directly grounding visual features with language models through RGB-based similarity scoring, significantly accelerating semantic processing. Nonetheless, the absence of explicit global layout-level guidance limits their ability to ensure efficient path planning. Works like HOVSG (werby23hovsg) construct open-vocabulary 3D scene topological graphs through feature clustering to enable navigation. However, the static topological way-points extracted from occupancy cannot ensure efficient path planning or manage unexpected local scene changes.
Overview
@@
We propose a hierarchical robotic navigation framework for operation in unexplored indoor environments, guided by natural language instructions $T$, images $I$, or 3D locations $P$. The framework overview is shown in Fig.2. The planning process is conducted on a dual-level map with global topology and local dense representation built with collected RGB-D sequences i,d{i,d}, where $i$ and $d$ denote RGB and depth observations, respectively. The scene representation consists of (1) an implicit neural function $F:\mathbb{R}^{3}\to\mathbb{R}^{n}$ maps 3D positions to vision-language embeddings; and (2) a topometric map $G=(V,E)$, where $V$ comprises region vertices $\mathbf{v}_{r}$ and entrance vertices connecting regions $\mathbf{v}_{e}$ and edges $E$ connecting these vertices.
@@
During planning, inputs $T/I$ are encoded into embeddings $\mathbf{E}$ using a pre-trained vision-language model. These embeddings are then queried in $F$ to identify the target position $p_{\text{end}}$ with the highest matching similarity. A global navigable path is computed between the robot’s current pose $p_{\text{start}}$ and $p_{\text{end}}$ using the topometric map $G$, generating a sparse sequence of waypoint vertices $V=\{v_{1},v_{2},\dots,v_{n}\}$. These waypoints are progressively refined during navigation in an imperative approach, with $W_{i,i+1}=\{w_{i}^{1},w_{i}^{2},\dots,w_{i}^{m}\}$ denoting the densified local waypoints between $v_{i}$ and $v_{i+1}$. Navigation proceeds incrementally: as the robot approaches $v_{i+1}$, the next segment $W_{i+1,i+2}$ is planned. The framework incorporates an LLM agent designed to integrate essential data streams for task execution.
Method
Hierarchical Scene Representation
@@
Recent works such as HOV-SG (werby23hovsg), CLIO (Maggio2024Clio), and Topo-Field (hou2024topofield) have demonstrated progress in representing scenes as topo-maps while retaining detailed content. Although topometric maps offer computational efficiency for downstream planning and navigation tasks, high-fidelity scene content from neural implicit representations enables semantic cue queries (hou2024topofield). Our approach represents 3D scenes combining a high-level topometric map and a detailed neural implicit representation.
@@
Given posed RGB-D sequences $\{i,d\}$, we train a neural implicit function $F$ (hou2024topofield) to match 3D points $p$ derived from depth and camera parameters with the encoded vision-language embedding $\mathcal{E}_{p}$ (clip; sentence-bert). A topometric map $G=(V,E)$ is then extracted with vertices $V$ including region vertices $\mathbf{v}_{r}$ and entrance vertices $\mathbf{v}_{e}$ and edges $E$ connecting them. The subsequent scene representation is denoted as:
$F$ $:\mathbb{R}^{3}\to\mathbb{R}^{n},$ (1)
@@
Details of neural field and topology construction pipeline can be referred to in (hou2024topofield).
Hierarchical Planning
@@
Global Topology Plan
To ensure globally efficient path planning for inter-room navigation, our approach utilizes a topological graph to generate a sparse waypoint set that the robotic must traverse to reach the destination, typically located at room entrances or connecting doors. The process begins by unifying multi-modal instructions (images, text, or 3D positions) into a destination 3D position. The text/image instruction $T/I$ is first encoded using CLIP (clip) and Sentence-BERT (sentence-bert) to generate the target embedding $\{C,S\}_{T/I}$. A point set $P^{*}$ is then sampled uniformly from the environment, and $F(P^{*})=\{(f_{v},f_{s})\}_{P^{*}}$, representing the vision and semantic embeddings of the corresponding 3D positions, is calculated. The cosine similarity between $\{C,S\}_{T/I}$ and $F(P^{*})$ determines the position $p_{\text{end}}\in P^{*}$ with the highest similarity to the instruction query (hou2024topofield). The topological graph $G$ is subsequently queried using the robot’s current position $p_{\text{start}}$ and $p_{\text{end}}$ to identify the shortest path connecting them. This results in a sequence of vertices $V=\{v_{1},v_{2},\dots,v_{n}\}$, where each $v_{i}$ represents a structural element, such as doors or corridors, that the robot must navigate through to reach the destination.
Local Ego-centric Plan
In the initial global planning phase, a sparse set of waypoints denoted as $V=\{v_{1},v_{2},\dots,v_{n}\}$, is established. These waypoints guide the robot’s inter-room navigation, ensuring efficient travel between distinct regions. To achieve intra-room movement and local navigation precision, dense waypoints, $W_{i,i+1}=\{w_{i}^{1},w_{i}^{2},\dots,w_{i}^{m}\}$, are computed between consecutive global waypoints $v_{i}$ and $v_{i+1}$. We realize local planning in an Imperative Learning (IL) approach with the target way-point $v_{i+1}$ in the global coordinates and current depth observation $d^{\prime}$ with pose $p^{\prime}$. First, $v_{i+1}$ is projected to the depth frame with the pose $p^{\prime}$. The depth input, encoded by a convolutional network (lecun1995convolutional) into a depth embedding $\mathcal{O_{d^{\prime}}}$, together with the way-point feature $\mathcal{P^{\prime}}$ of $p^{\prime}$ is input into a planning network (yang2023iplanner) that generates a set of local key points $W_{i,i+1}$ from
$\tau=\mathcal{F}(\mathcal{O}_{d^{\prime}},\mathcal{P^{\prime}}),$ (2)
@@
$p_{\text{start}}\to p_{\text{end}}$ $:V=\{v_{1},v_{2},\dots,v_{n}\},$ (3)
@@
LLM-powered Navigation Agent
Figure 4: Data flow example of LLM Agent. Instructed by the user input, the agent conducts prompts, applies optional actions, and gets feedback from interaction.
@@
CoW (2022) (gadre2022cow) 0.061 0.039 0.018 0.011 - -
ZSON (2022) (majumdar2022zson) 0.255 0.126 0.106 0.049 0.146 0.073
L3MVN (2023) (yu2023l3mvn) 0.352 0.165 - - - -
ESC (2023) (zhou2023esc) 0.392 0.223 0.065 0.037 - -
VLFM (2024) (yokoyama2024vlfm) 0.364 0.175 - - - -
PixNav (2024) (cai2024pixnav) 0.379 0.205 - - - -
HOV-SG (2024) (werby23hovsg) 0.404 0.236 - - - -
PSL (2024) (psl) 0.424 0.192 0.165 0.075 0.230 0.114
LOG-Nav(Ours) 0.856 0.797 0.786 0.701 0.673 0.582
Table 1: Quantitative comparison of object navigation and instance navigation tasks on MP3D dataset. For object navigation, text instruction is used as input. For instance navigation, text-prompted and image-prompted instructions are separately validated. The SR and SPL are employed as metrics.
@@
Table 2: Quantitative object navigation evaluation in real-world scene. Each navigation task is evaluated 10 times from a similar starting position to the same target. Obstacles means the number of obstacles we set after robotic exploration where robotics have to bypass to approach targets. Local/Global-Plan Retries means the number of replanning called by the agent during the 10 times of navigation. Path Length is the average path length robotics travel in 10 times navigation. Time cost is the average cost from starting points to destinations, where failure cases are not counted. SR is the navigation success rate.
@@
$\{\mathbf{B},\mathbf{I},\mathbf{M},\mathbf{P},\mathbf{T},\mathbf{S},\mathbf{O},\mathbf{A}\}$ (4)
where $\mathbf{B}$ represents the task background information, structured as “You are an embodied robotic …”. $\mathbf{I}$ denotes the user’s instruction, which can be text, an image, or a 3D position. $\mathbf{M}$ is the hierarchical map of the environment, comprising a learned implicit function $F$ and a topological graph $G$ in JSON format. $\mathbf{P}$ includes the robot’s historical planning data. $\mathbf{T}$ refers to the robotic travel trajectory. $\mathbf{S}$ indicates the robot’s current status, such as its location and whether the planned path remains viable. $\mathbf{O}$ represents any optional skills the robot may use. $\mathbf{A}$ is the action decision.
@@
  • Planning. Based on the scene representation $(F,G)$ and user instructions, the agent conducts planning described in Section Hierarchical Planning to fulfill the task.